# Hybrid quantum lattice model: Polaritons, photons, and spin waves propagation

### v2

@@ Abstract
Controlling the propagation of quantum excitations in low-dimensional systems is pivotal for quantum technologies, including communication networks and quantum simulators. We propose a one-dimensional hybrid quantum lattice model, where each lattice unit integrates a single-mode resonator that interacts with a two-level system (TLS), featuring direct coupling between adjacent TLSs. This configuration enables the coherent propagation of polaritons, spin waves, and photons, depending on the interplay between light-matter coupling and qubit-qubit interactions. Employing the time-evolving block decimation (TEBD) algorithm, we simulate the dynamics of various excitation configurations and analyze their transport characteristics using local observables. Our analysis reveals the importance of matching impedance and resonance conditions via system parameters for the propagation of different types of excitations or swapping the nature of excitations along the hybrid lattice. These findings offer insights into designing controllable quantum links and single-excitation swaps in low-dimensional quantum systems.
@@ I Introduction
The precise control of excitation propagation in quantum systems is a fundamental requirement for quantum communication networks [1, 2], quantum simulators [3, 4], energy transfer devices [5, 6], quantum state transfer [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], and distributed quantum computing [19, 20, 21, 22, 23, 24, 25]. In particular, one-dimensional (1D) quantum lattices provide an ideal platform for exploring the dynamics of different excitation types, including photons, polaritons, and spin waves. Advancements in cavity QED [26, 27], spin chain systems [28], trapped ions [29, 30], and superconducting circuits [31, 32, 33] have enabled the realization of hybrid models where light-matter interactions and qubit-qubit couplings can be engineered to tailor excitation dynamics.
A scenario where different types of excitations appear is the Jaynes-Cummings-Hubbard (JCH) model [34, 35, 36], a paradigmatic model for studying quantum phase transitions. Their static and dynamical properties have been a subject of great interest [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Recently, the JCH model has been proposed for quantum state transfer in short-range communication lattice [52, 53]. One may wonder what other scenarios, physically implementable in quantum simulators [3, 4], may also support the propagation of different types of excitations on-chip.
This work addresses the previous concern by proposing a 1D hybrid quantum lattice model where each lattice unit integrates a single-mode resonator that interacts with a two-level system. This configuration enables the propagation of distinct excitation types: polaritons, spin waves, and photons. Additionally, the proposed model enables the swapping of excitations based on the relative strengths of the light-matter coupling and qubit-qubit interactions. By tuning these parameters, we focus on controlling the transport properties of each excitation type within a unified framework.
To analyze the system dynamics, we employ the time-evolving block decimation (TEBD) algorithm [54, 55, 56], which allows efficient simulation of nonequilibrium many-body systems. Our simulations reveal transport characteristics for each excitation type and highlight the role of system parameters in modulating propagation behavior. These results provide a foundation for designing controllable quantum links [17, 18, 57] and single-excitation swapping, with potential applications in quantum information processing, energy transfer, and the simulation of strongly correlated light-matter systems.
@@ II Model
The hybrid quantum lattice consists of light-matter units, each comprising a single-mode resonator interacting with a single two-level system (TLS), as shown schematically in Fig. 1(a). In addition to the light-matter interaction, the TLSs interact directly with their nearest neighbors. An activation qubit (emitter) is placed at the first site of the hybrid quantum lattice, and its primary role is to inject a single excitation into the hybrid lattice. The activation qubit is directly coupled to the TLS of the first unit in the hybrid lattice, located at site $j=2$. The hybrid quantum lattice supports the propagation of three types of single excitations: photons, polaritons, and spin waves. The proposed model is inspired by state-of-the-art circuit QED experiments [31, 58, 59, 60] as schematically shown in Fig. 1(b). Here, a set of transmons, represented by crosses, are capacitively coupled in a linear array. Each light-orange transmon is capacitively coupled to a single-mode resonator (wavy lines). We stress that in Refs.[31, 58, 59, 60], the resonators were used for measuring the transmon state. However, there are no fundamental physical limitations preventing the transmon-resonator system from strongly interacting and forming polaritons.
@@
where $\hat{a}_{j}^{{\dagger}}$ ($\hat{a}_{j}$) is the creation (annihilation) operator of a single bosonic excitation in the $j$-th resonator with frequency $\omega$, and $\hat{\sigma}_{j}^{+}$ and $\hat{\sigma}^{-}_{j}$ are the raising and lowering operators of the $j$-th TLS in the eigenbasis $\{|\uparrow_{j}\rangle,|\downarrow_{j}\rangle\}$. Additionally, $\omega_{\rm{A}}$ denotes the frequency of the activation qubit, $\omega_{0}$ the transition frequency of the TLSs in the hybrid lattice, $g$ the light-matter coupling strength, $\lambda$ the coupling strength between the activation qubit and the TLS in the first unit of the hybrid lattice, and $v$ the coupling strength between nearest-neighbor TLSs.
To analyze the propagation of different types of single excitations, it is convenient to rewrite the full Hamiltonian in the basis of hybrid light-matter states, i.e., the upper (+)(+)( + ) and lower ()(-)( - ) polaritonic basis states
@@
$E^{\alpha}_{n}/\hbar=n\omega+\frac{\Delta}{2}\pm\frac{\Delta}{2}\sqrt{1+\frac{4ng^{2}}{\Delta^{2}}}\,.$ (6)
@@
III Multiple excitations propagation
In this section, we expose the propagation mechanism of different types of excitations, namely, polaritons, spin waves, and photons, within a unified platform: the hybrid quantum lattice. The propagation dynamics of each excitation type depend on tuning key parameters: the light–matter coupling strength, the activation qubit frequency, and the coupling between neighboring TLSs. As will be shown later, by tuning these parameters and suppressing the excitation propagation through the upper polaritonic branch, we can select the type of information carrier propagating in the system. We focus on two distinct cases of light-matter interaction: the resonant and dispersive regimes. In the resonant regime, where $\Delta=\omega_{0}-\omega=0$, polaritons dominate the dynamics. In contrast, in the dispersive regime, where $|\Delta|\gg g$, the single photon and spin-wave excitation become the dominant information carriers.
We perform numerical simulations using the TEBD algorithm [54, 55, 56]. We numerically approximate the evolution operator using the Suzuki-Trotter decomposition [64]
@@ III Multiple excitations propagation
where $F$ and $G$ contain the sum of the Hamiltonian terms for the even and odd bonds [54], respectively. In particular, we use the second-order ($n=2$) approximant, where the expansion parameters of Eq. (10) are $k=2$, $c_{1}=1/2$, $d_{1}=1$, $c_{2}=1/2$, $d_{2}=0$. Besides, we consider a time-step $\tau=10^{-3}v^{-1}$ for spin and polariton propagation, and $\tau=5\times 10^{-3}v^{-1}$ for photon propagation, the total system size of $L=26$ sites, the accessible number states in each resonator is $n_{\rm max}=2$, and the maximum bond dimension is $\chi_{\rm max}=4$. Within the simulated time, the latter ensures a truncation error $\epsilon_{\rm trunc}\lesssim 10^{-8}$, see Appendix B.
@@
[…]
@@ III.1 Resonant regime: polariton propagation
Figure 2: Polariton-wave propagation: (a) Time evolution of the average excitation number in the activation qubit. (b) Time evolution of the average excitation number in the $j$th TLS of the hybrid lattice, for $2\leq j\leq 26$. (c) Average photon number as a function of time. (d) Average polariton number as a function of time and the site index $j$. The parameters are $\Delta=0$ with $\omega=\omega_{0}=1$, $v=0.01\,\omega$, $g=4\,v$, $\omega_{\rm{A}}=\omega_{pol}=\omega-g$, $\lambda=-v/\sqrt{2}$, $\tau=10^{-3}v^{-1}$, and $L=26$.
@@
Tuning the lattice parameters to satisfy the three previously described conditions enables the propagation of a single polariton through the hybrid quantum lattice. Figure 2 shows the mechanism of single-excitation generation, its release, and subsequent polariton propagation. Figure 2(a) displays the behavior of the activation qubit, where the average excitation number $\langle\hat{\sigma}^{+}_{1}\hat{\sigma}^{-}_{1}\rangle=1$ at $t=0$, and eventually decays to 0, injecting a single excitation into the hybrid system. Once released, the excitation propagates to the right as a single polariton. Figures 2(b) and 2(c) show the average excitation number of the $j$th TLS and the average photon number in each resonator, respectively. Since the polariton number operator is defined as $\hat{\mathcal{N}}_{pol,j}=\hat{\sigma}^{+}_{j}\hat{\sigma}^{-}_{j}+\hat{n}_{j}$, these two figures separately depict the respective contributions of spin and photon components to the polariton dynamics. In Fig. 2(d), we show the average polariton number $\langle\hat{\mathcal{N}}_{pol,j}\rangle$ as a function of time and site index $j$. The polaritonic excitation appears as a dispersive pulse, with a decreasing amplitude and increasing width over time.
@@ III.2 Dispersive regime: photon propagation
Figure 3: Photon-wave propagation: (a) Time evolution of the average excitation number in the activation qubit. (b) Time evolution of the average excitation number in the $j$th TLS of the hybrid lattice, for $2\leq j\leq 26$. (c) Average photon number as a function of time. (d) Average photon number as a function of time and the site index $j$. The parameters are $\omega=10$, $\omega_{0}=1$, $\Delta=\omega_{0}-\omega$, $g=0.08\,\omega$, $\omega_{\rm A}=\omega-g^{2}/\Delta$, $v=0.05\,\omega$, $\lambda=-v\,\sin{\left(\theta_{1}/2\right)}$ with $\theta_{1}=\arctan{(2g/\Delta)}$, $\tau=5\times 10^{-3}v^{-1}$, and $L=26$.
In the dispersive regime ($|\Delta|\gg g$), single-photon propagation through the hybrid lattice is possible even when the average excitation number of the $j$th TLS remains approximately zero. In this regime, the photon transport is mediated by the hybrid light-matter excitations, namely the polaritons. For photon propagation to occur, the following conditions must be satisfied:
@@
    $E^{-}_{n}/\hbar\simeq\omega+\frac{\Delta}{2}-\frac{\Delta}{2}\left(1+\frac{\chi(n)}{2}\right)\,=\left(\omega-\frac{g^{2}}{\Delta}\right)n\,.$ (19)
@@
Figure 4: Spin-wave propagation: (a) Time evolution of the average excitation number in the activation qubit. (b) Time evolution of the average excitation number in the $j$th TLS of the hybrid lattice, for $2\leq j\leq 26$. (c) Average photon number as a function of time. (d) Average excitation number in the $j$th TLS as a function of time and the site index $j$. The parameters are $\omega=1$, $\omega_{0}=10$, $\Delta=\omega_{0}-\omega$, $\omega_{\rm{A}}=\omega_{0}$, $g=0.05\,\omega$, $v=0.02\,\omega$, $\lambda=v$, $\tau=10^{-3}v^{-1}$, and $L=26$.
III.3 Dispersive regime: Spin wave propagation
In the dispersive regime ($|\Delta|\gg g$), a second case of interest is the propagation of spin-wave excitations. This regime allows for excitation transport through the TLS sublattice, while the photonic component remains largely unexcited. The following conditions must be satisfied:

### v3

# Single excitation swap in a modified Jaynes-Cummings-Hubbard lattice

@@ Abstract
Controlling the transport and nature of quantum excitations in low-dimensional systems is a key requirement for scalable quantum devices, including communication networks and quantum simulators. We propose a one-dimensional hybrid quantum lattice model, in which each lattice unit integrates a single-mode resonator that interacts with a two-level system (TLS), featuring direct coupling between adjacent TLSs. This configuration enables the coherent propagation of excitations with tunable atomic, photonic, or polaritonic character. Beyond conventional single-excitation transport, we demonstrate that appropriate impedance-matching and resonance conditions allow for the controlled swapping of excitation type as the excitation propagates along the lattice. We analyze the resulting dynamics using local observables and pairwise concurrence to track both transport and quantum correlations. Our results establish a minimal platform for controlled single-excitation conversion, with direct relevance to hybrid quantum networks, on-chip quantum interconnects, and engineered quantum simulators.
@@ I Introduction
The ability to control the propagation of quantum excitations in engineered many-body systems is a central requirement for quantum communication networks [1, 2, 3], quantum simulators [4, 5], quantum state transfer protocols [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], distributed quantum computing architectures [17, 18, 19, 20, 21, 22, 23], and energy transfer devices [24, 25]. One-dimensional (1D) quantum lattices provide an especially attractive setting for these tasks, as they enable controlled transport of single or few excitations while remaining amenable to both analytical methods and numerical simulations. In such lattices, excitations may take different physical forms, including spin waves [26, 27], bosons [28, 29], hybrid light-matter excitations [30, 31, 32, 33], and photons [34, 35, 36]. Nowadays, 1D quantum lattices can be implemented in circuit quantum electrodynamics [37, 38, 39], trapped ions [40, 41, 42], and Rydberg atoms [43], enabling the realization of the hybrid architectures in which light-matter coupling and inter-qubit interactions can be engineered with high precision.
A paradigmatic framework capturing the dynamics of the photonic, atomic, or hybrid excitations is provided by the Jaynes-Cummings-Hubbard (JCH) model [30, 31, 32, 44, 45]. Originally introduced in the context of quantum phase transitions, the JCH model has since been explored in a wide range of static and dynamical regimes [46, 33, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. It has also been proposed as a platform for quantum state transfer in both long-range [60] and short-range lattice geometries [61, 62]. These works have established the feasibility of coherent single-excitation transport in hybrid light-matter systems. However, the ability to dynamically swap the character of an excitation during its propagation, for example, an atomic excitation into a polaritonic one and vice versa, remains unexplored, despite its potential relevance for hybrid quantum networks and on-chip quantum interconnects.
In this work, we address this gap by proposing a one-dimensional hybrid lattice in which each unit consists of a single-mode resonator locally coupled to a two-level system, while neighbouring units interact directly through their TLSs, see the schematic of Fig.1. This architecture is a modified JCH model that naturally supports the propagation of polaritonic and purely atomic excitations. However, purely photonic excitations require adding an ancillary emitter that mediates the conversion between excitation types. We prove that, under suitable impedance-matching and resonance conditions derived in the hybrid light-matter basis, the system enables a controlled swapping of excitation type at a local interface within the lattice. To establish this mechanism, we revisit how a single atomic excitation, initially prepared in the emitter, can be injected into the lattice and converted into propagating purely atomic, purely photonic, or polaritonic excitations by parameter tuning. This analysis provides the necessary framework to identify the conditions under which coherent excitation-type swapping occurs dynamically during propagation. The dynamical swapping and conversion mechanisms are studied by using the average of local observables and entanglement via pairwise concurrences [63].
The proposed mechanism relies on parameter regimes that are accessible in circuit-QED architectures [64], making it experimentally feasible with current technology. Our results establish a minimal and versatile platform for controlled single-excitation conversion and swapping, with applications to the design of controllable quantum links [15, 16, 65], hybrid quantum networks, and engineered quantum simulators of light-matter systems.
@@ II Model
The hybrid quantum lattice consists of light-matter units, each comprising a single-mode resonator interacting with a single two-level system, as shown schematically in Fig. 1(a). In addition to the light-matter interaction, the TLSs interact directly with their nearest neighbors. An activation qubit (emitter) is placed at the first site of the hybrid quantum lattice, and its primary role is to inject a single excitation into the hybrid lattice. The activation qubit is directly coupled to the TLS of the first unit in the hybrid lattice, located at site $j=2$. The hybrid quantum lattice supports the propagation of three types of single excitations: polaritons, purely atomic excitations, and purely photonic excitations. The proposed model is inspired by state-of-the-art circuit QED experiments [37, 66, 67, 68] as schematically shown in Fig. 1(b). Here, a set of transmons, represented by crosses, are capacitively coupled in a linear array. Each light-orange transmon is capacitively coupled to a single-mode resonator (wavy lines). We stress that in Refs.[37, 66, 67, 68], the resonators were used for measuring the transmon state. However, there are no fundamental physical limitations preventing the transmon-resonator system from strongly interacting and forming polaritons.
@@
where $\hat{a}_{j}^{{\dagger}}$ ($\hat{a}_{j}$) is the creation (annihilation) operator of a single bosonic excitation in the $j$-th resonator with frequency $\omega$, and $\hat{\sigma}_{j}^{+}$ and $\hat{\sigma}^{-}_{j}$ are the raising and lowering operators of the $j$-th TLS in the eigenbasis $\{|\uparrow_{j}\rangle,|\downarrow_{j}\rangle\}$. Additionally, $\omega_{\rm{A}}$ denotes the frequency of the activation qubit, $\omega_{0}$ the transition frequency of the TLSs in the hybrid lattice, $g$ the light-matter coupling strength, $\lambda$ the coupling strength between the activation qubit and the TLS in the first unit of the hybrid lattice, and $v$ the coupling strength between nearest-neighbor TLSs. Along this work, we deal with a single excitation propagation. In this case, the Hamiltonian of our model can be mapped to an effective single-particle problem and is therefore formally related to the JCH systems studied in Ref. [33].
To determine impedance-matching conditions and analyze the propagation of different types of single excitations, it is convenient to rewrite the full Hamiltonian in the basis of hybrid light-matter states. The methodology for performing the basis transformation has been reported previously for the Jaynes-Cummings-Hubbard model in Refs. [32, 46]. Then, we denote the upper $(+)$ and lower ()(-) polaritonic basis states are
@@
$E^{\pm}_{n}/\hbar=n\omega+\frac{\Delta}{2}\pm\frac{\Delta}{2}\sqrt{1+\frac{4ng^{2}}{\Delta^{2}}}\,.$ (6)
@@
III Single-excitation propagation in atomic, photonic, and polaritonic channels
Previous studies have examined the three propagation regimes of the JCH model in different contexts. The atomic propagation regime was analysed by Ogden et al. in Ref. [44], using a minimal two-site model and a description in terms of delocalized field modes, where atomic excitation transfer can occur with negligible field population. A generalization of this mechanism to larger arrays was later presented by Chakrabarti et al. in Ref. [45], who used delocalized photonic and atomic modes to describe predominantly atomic or photonic propagation under appropriate detuning conditions. Photonic and hybrid excitation transport has also been studied in Ref. [33], where the atomic signal corresponds to the atomic component of a hybrid excitation. In addition, Ref. [60] investigated atomic-state transfer mediated indirectly through hybrid excitation in a staggered JCH lattice, where the excitation acquires a non-negligible photonic component in the bulk during its propagation.
In this section, we present the propagation of different types of excitations, that is, polaritons, atomic waves, and photons in our model. This analysis provides a framework for proposing the main contribution of our work, namely, the excitation-swap mechanism. Our methodology is not focused on delocalized collective fields, but rather derives an impedance-matching and resonance condition formulated using the Hamiltonian transformation to the polaritonic basis, along with a large detuning of the local fields. In our model, the propagation dynamics of each excitation type strongly depend on suitable impedance-matching and resonance conditions. As will be shown later, by satisfying these conditions and suppressing the excitation propagation through the upper polaritonic branch, we can select the type of information carrier propagating in the system. We focus on two distinct cases of light-matter interaction: the resonant and dispersive regimes. In the resonant regime, where $\Delta=\omega_{0}-\omega=0$, polaritons dominate the dynamics. In contrast, in the dispersive regime, where $|\Delta|\gg g$, the single photon and atomic-wave excitation become the dominant information carriers.
We perform numerical simulations using the TEBD algorithm [72, 73, 74]. We approximate the evolution operator using the Suzuki-Trotter decomposition [75]
@@ III Single-excitation propagation in atomic, photonic, and polaritonic channels
where $F$ and $G$ contain the sum of the Hamiltonian terms for the even and odd bonds [72], respectively. In particular, we use the second-order ($n=2$) approximant, where the expansion parameters of Eq. (10) are $k=2$, $c_{1}=1/2$, $d_{1}=1$, $c_{2}=1/2$, $d_{2}=0$. Besides, we consider a time-step $\tau=10^{-3}v^{-1}$ for atomic and polariton propagation, and $\tau=5\times 10^{-3}v^{-1}$ for photon propagation, the total system size of $L=26$ sites, the accessible number states in each resonator is $n_{\rm max}=2$, and the maximum bond dimension is $\chi_{\rm max}=4$. Within the simulated time, the latter ensures a truncation error $\epsilon_{\rm trunc}\lesssim 10^{-8}$, see Appendix B. We emphasize that the system evolves within the single-excitation subspace, allowing for an equivalent description in terms of a finite-dimensional Hamiltonian. Nevertheless, we employ the TEBD algorithm as a flexible and scalable numerical approach that can readily be extended to higher numbers of excitations, involving larger local Hilbert spaces. In this sense, TEBD provides a consistent framework for future studies of multi-excitation dynamics.
@@
$|\Psi_{0}\rangle=|\uparrow_{1}\,\rangle\otimes|\downarrow,\,0\,\rangle_{2}\otimes|\downarrow,\,0\,\rangle_{3}\otimes\dots\otimes|\downarrow,\,0\,\rangle_{L}\,.$ (11)
@@ III.1 Resonant regime: polariton propagation
Figure 2: Polariton-wave propagation: (a) Time evolution of the average excitation number in the activation qubit. (b) Time evolution of the average excitation number in the $j$th TLS of the hybrid lattice, for $2\leq j\leq 26$. (c) Average photon number as a function of time. (d) Average polariton number as a function of time and the site index $j$. The parameters are $\Delta=0$ with $\omega=\omega_{0}=1$, $v=0.01\,\omega$, g=4vg=4\,v, $\omega_{\rm{A}}=\omega_{pol}=\omega-g$, $\lambda=-v/\sqrt{2}$, $\tau=10^{-3}v^{-1}$, $\chi_{\rm max}=4$, and $L=26$.
@@
Tuning the lattice parameters to satisfy the three previously described conditions enables the propagation of a single polariton through the hybrid quantum lattice. Figure 2 shows the mechanism of single-excitation generation, its release, and subsequent polariton propagation. Figure 2(a) displays the behavior of the activation qubit, where the average excitation number $\langle\hat{\sigma}^{+}_{1}\hat{\sigma}^{-}_{1}\rangle=1$ at $t=0$, and eventually decays to 0, injecting a single excitation into the hybrid system. Once released, the excitation propagates to the right as a single polariton. Figures 2(b) and 2(c) show the average excitation number of the $j$th TLS and the average photon number in each resonator, respectively. Since the polariton number operator is defined as $\hat{\mathcal{N}}_{pol,j}=\hat{\sigma}^{+}_{j}\hat{\sigma}^{-}_{j}+\hat{n}_{j}$, these two figures separately depict the respective contributions of atomic and photon components to the polariton dynamics. In Fig. 2(d), we show the average polariton number $\langle\hat{\mathcal{N}}_{pol,j}\rangle$ as a function of time and site index $j$. The polaritonic excitation appears as a dispersive pulse, with a decreasing amplitude and increasing width over time.
@@ III.2 Dispersive regime: photon propagation
Figure 3: Photon-wave propagation: (a) Time evolution of the average excitation number in the activation qubit. (b) Time evolution of the average excitation number in the $j$th TLS of the hybrid lattice, for $2\leq j\leq 26$. (c) Average photon number as a function of time. (d) Average photon number as a function of time and the site index $j$. The parameters are $\omega=10$, $\omega_{0}=1$, $\Delta=\omega_{0}-\omega$, $g=0.08\,\omega$, $\omega_{\rm A}=\omega-g^{2}/\Delta$, $v=0.05\,\omega$, $\lambda=-v\,\sin{\left(\theta_{1}/2\right)}$ with $\theta_{1}=\arctan{(2g/\Delta)}$, $\tau=5\times 10^{-3}v^{-1}$, $\chi_{\rm max}=4$, and $L=26$.
In the dispersive regime ($|\Delta|\gg g$), purely photonic transport may be mediated by the hybrid light-matter excitations. For photon propagation to occur, the following conditions must be satisfied:
@@
    $E^{-}_{n}/\hbar\simeq n\omega+\frac{\Delta}{2}-\frac{\Delta}{2}\left(1+\frac{\chi(n)}{2}\right)\,=\left(\omega-\frac{g^{2}}{\Delta}\right)n\,.$ (19)
@@
Figure 4: Atomic-wave propagation: (a) Time evolution of the average excitation number in the activation qubit. (b) Time evolution of the average excitation number in the $j$th TLS of the hybrid lattice, for $2\leq j\leq 26$. (c) Average photon number as a function of time. (d) Average excitation number in the $j$th TLS as a function of time and the site index $j$. The parameters are $\omega=1$, $\omega_{0}=10$, $\Delta=\omega_{0}-\omega$, $\omega_{\rm{A}}=\omega_{0}$, $g=0.05\,\omega$, $v=0.02\,\omega$, $\lambda=v$, $\tau=10^{-3}v^{-1}$, $\chi_{\rm max}=4$, and $L=26$.
III.3 Dispersive regime: Atomic wave propagation
In the dispersive regime ($|\Delta|\gg g$), a second case of interest is the propagation of atomic-wave excitations. This regime allows for excitation transport through the TLS sublattice, while the photonic component remains largely unexcited. The following conditions must be satisfied: